OpenStudy (anonymous):

the midpoint of ST is (5,-8) find the coordinates of point S. T(10,18)

OpenStudy (anonymous):

\[\left( \frac{ x1+x2 }{ 2 } ,\frac{ y1+y2 }{ 2 }\right)\]